OpenStudy (anonymous):

An object weighing 49N is pushed across a floor by a force of 12 N. What is the acceleration of the object?

OpenStudy (anonymous):

omg gosh yes finally someone to help me

OpenStudy (anonymous):

Assuming the floor is frictionless?

OpenStudy (anonymous):

Without any slope, the acceleration is equal to the force of push divided by the objects mass.

OpenStudy (anonymous):

what does that even mean

OpenStudy (anonymous):

When you push an object to the right, it moves to the right, so the only forces that are acting upon it to make it move, are the ones that are parallel to the horizon. In this case, it's the force that pushes the object at 12 N. Also, Newton says that F = ma. So you can divide by the mass to get F/m = a. Where the mass is given by the objects weight, and the force is the one that's actually making it move.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

where would i get the mass from

OpenStudy (anonymous):

The weight of an object is equal to its mass times the earth gravitational acceleration constant "g", usually given as 9.81 m/s. w = m * g The problem states that the objects weight is 49 N.

OpenStudy (anonymous):

is it 2.40

OpenStudy (anonymous):

Technically the g in this case would be called the gravitational field strength not acceleration due to gravity because you are not in freefall. \[g = 9.8 \frac{N}{kg}\] However the numbers would be the same

OpenStudy (anonymous):

Yes 2.4 m/s^2 is correct for the acceleration

OpenStudy (anonymous):

yay thank u